# Certificate-Sensitive Subset Sum: Realizing Instance Complexity

### v4

@@ Appendix E Controlled Aliasing: Correctness and Implementation
Let $S^{\prime}\subseteq\mathbb{Z}_{>0}$ be a set of $k$ positive integers, and let $x_{0},x_{1}\in S^{\prime}$ be a designated aliasing pair with $x_{0}\neq x_{1}$. Suppose we apply the aliasing rule during subset sum enumeration: whenever $x_{1}$ appears in a subset, its value is treated as $x_{0}$. Then the total number of distinct subset sums generated is at most $\frac{3}{4}\cdot 2^{k}$.
@@
Partition the $2^{k}$ subsets of $S^{\prime}$ according to whether they include neither, one, or both of $x_{0}$ and $x_{1}$. Let $P\subseteq S^{\prime}\setminus\{x_{0},x_{1}\}$ be an arbitrary subset of the remaining elements.
There are four corresponding supersets of $P$ formed by optionally adding $x_{0}$ and/or $x_{1}$:
  • $P$
  • $P\cup\{x_{0}\}$
  • $P\cup\{x_{1}\}$
  • $P\cup\{x_{0},x_{1}\}$
Under normal enumeration, each contributes a distinct sum. Under the aliasing rule, $x_{1}$ is treated as $x_{0}$, so the last three all contribute sums of the form $\sigma(P)+x_{0}$, except the last, which contributes $\sigma(P)+2x_{0}$. Thus, the four subsets yield only three unique sums. Since there are $2^{k-2}$ such base subsets $P$, the total number of unique subset sums is at most $3\cdot 2^{k-2}=\frac{3}{4}\cdot 2^{k}$. ∎
Core Procedures
Alias State Function.
@@
2:  return $2\cdot b[i_{1}]+b[i_{0}]$
Insert into Memo Table.
@@
2:  $a\leftarrow\textsc{AliasState}(b,i_{0},i_{1})$
[…]
Corrected Sum.
[…]
2:  if $a=2$ or $a=3$ then
3:   return $\sigma-x_{0}+x_{1}$
4:  else
5:   return $\sigma$
Aliased Value (Inverse).
[…]
2:  if $a=2$ or $a=3$ then
3:   return $\sigma_{real}+x_{0}-x_{1}$
4:  else
5:   return $\sigma_{real}$
@@
To ensure correctness during merging, the algorithm performs a 16-way compensatory check across alias states of both halves. For each left-side sum, it computes the required right-side partner under all alias state interpretations.
@@
[…]
2:  for all $\sigma_{L}\in\texttt{MemoLeft}$ do
3:   for $a_{L}=0$ to 333 do
4:     $b_{L}\leftarrow\texttt{MemoLeft}[\sigma_{L}][a_{L}]$
@@
6:     $s_{L}\leftarrow\textsc{CorrectedSum}(\sigma_{L},a_{L},x_{0},x_{1})$
@@
8:     for $a_{R}=0$ to 333 do
9:      $\sigma_{R}\leftarrow\textsc{AliasedValue}(s_{R},a_{R},y_{0},y_{1})$
10:      if $\sigma_{R}\in\texttt{MemoRight}$ then
11:        $b_{R}\leftarrow\texttt{MemoRight}[\sigma_{R}][a_{R}]$
@@
13:         return $b_{L}\,\|\,b_{R}$
@@
  • Each reachable sum stores up to 4 bitmasks.
  • All routines operate in constant time per alias lane.
  • Total time and space remain:
[…]
  • Merge routine performs $O(U)$ lookups with constant overhead.
Conclusion
This 4-lane structure solves the canonical aliasing problem completely. It preserves semantic distinctions between aliased subsets, maintains lexicographic minimality in each lane, and supports full witness reconstruction without loss of correctness or efficiency. The Controlled Aliasing mechanism now guarantees correctness for both decision and construction variants of Subset–Sum on all inputs.
Appendix F Note: Decision vs. Enumeration Hardness
[…]
  • Perform a prefix-guided traversal of the subset tree using $\mathcal{A}$ as an oracle.
  • At each node, extend the prefix by one bit (include or exclude the next element).
  • Query whether any extension leads to a feasible sum in the instance.
  • Prune infeasible branches, recurse on feasible ones, and memoize discovered sums.
[…]

### v6

@@ Appendix C Compositional Sum Lemmas and Complement Check
[…]
9:function TestDirectOnRight($s$) $\triangleright$ seek $B$ with $\sigma(B)=t-s$
10:  for $\chi_{R}\in\{0,1\}$ do
11:   $\tilde{\sigma}_{R}\leftarrow\textsc{AliasedValue}\big{(}t-s,\chi_{R},u_{0},u_{1}\big{)}$
[…]
13:     return true
14:  return false
15:function TestRightComplement($s$) $\triangleright$ seek $B$ with $\sigma(\ell_{b^{\prime}}\!\setminus\!B)=t-s$
16:  $g\leftarrow\mathrm{Sum}_{b^{\prime}}-(t-s)$ $\triangleright$ goal real sum for $B$
17:  for $\chi_{R}\in\{0,1\}$ do
18:   $\tilde{\sigma}_{R}\leftarrow\textsc{AliasedValue}\big{(}g,\chi_{R},u_{0},u_{1}\big{)}$
[…]
20:     return true
21:  return false Intra-half direct/self-complement:
22:if $s_{\mathrm{real}}=t$ or $\mathrm{Sum}_{b}-s_{\mathrm{real}}=t$ then
23:  Output: direct solution found
24:  return Cross-half patterns:
25:if TestDirectOnRight($s_{\mathrm{real}}$then
26:  Output: direct across splits; return
27:if TestRightComplement($s_{\mathrm{real}}$then
28:  Output: mixed (right-complement); return
[…]
30:if TestDirectOnRight($s_{\mathrm{comp}}$then
31:  Output: mixed (left-complement); return
32:if TestRightComplement($s_{\mathrm{comp}}$then
33:  Output: double complement; return
Appendix D Experimental Details
We provide additional data and methodology for the experiments summarized in Section 7. Our aim is not to claim new empirical bests, but to validate the central prediction of our theory: for fixed $n$, the wall-time of IC-SubsetSum closely tracks the size of the constructive certificate $U=|\Sigma(S)|$.
D.1 Setup and Methodology
Implementation.
All experiments use a single-threaded, 64-bit C++17 implementation compiled with -O3 and no target-specific vectorization. Wall-clock time is measured with high-resolution timers; each configuration is run for multiple independent seeds and we report the median together with the interquartile range (IQR). No parallelism or GPU acceleration is used.
Metrics.
For each instance we record:
  • $U=|\Sigma(S)|$ (or per-half $U_{i}=|\Sigma(\ell_{i})|$ when using meet-in-the-middle),
  • the collision entropy $H_{c}(S)=n-\log_{2}U$,
  • total number of extension attempts and collision resolutions, and
  • wall-clock time $T$ for full constructive-certificate enumeration.
When $U$ is small enough to store exactly, we count it via exact hashing. For larger $n$ where exact counting may be infeasible under memory limits, we estimate $U$ using a standard streaming distinct-counter (with sub-percent relative error at the configured memory budget); in those cases, only trend lines are reported. All qualitative trends are consistent with exact counts on smaller instances.
Workloads.
We vary three structure “knobs” known to affect collisions:
  1. a)
    Numeric density: draw $a_{i}\sim\mathrm{Unif}(\{1,\dots,2^{w}\})$ for several word sizes $w$,
  2. b)
    Duplicates: start from a base multiset and inject $d\in\{0,2,4,\dots\}$ extra copies of randomly chosen elements, and
  3. c)
    Additive progressions: splice arithmetic-progressions of fixed length into otherwise uniform data.
Unless noted otherwise, we sort inputs once (for cache locality) but the algorithm does not rely on ordering for correctness.
D.2 Extremely Dense Instances ($n=100$)
Here we stress-test the density knob by fixing $n=100$ and drawing elements from very small ranges $1,2^{w}]$ with $w\in\{12,16,20,24\}$. As $w$ decreases, collisions become pervasive and $U$ collapses far below the collision-free ceiling. In meet-in-the-middle terms, the per-half certificate sizes $U_{0},U_{1}$ drop by orders of magnitude relative to $2^{n/2}$.
Observation. Across densities, we observe a monotone relationship between $T$ and $U$: for fixed $n$, the median wall-time scales linearly with $U$ (up to polynomial factors in $n$), in line with Theorem 1 (deterministic $O(U\cdot n^{2})$) and Theorem 2 (expected $O(U\cdot n)$). In particular, once $U\ll 2^{n/2}$ due to density-induced collisions, runtime drops commensurately.
Practical note.
For $n=100$ and large $w$, exact per-half certificates may exceed memory on commodity hardware; in those regimes we report consistent trends using streaming distinct-count estimates and terminate runs upon reaching a fixed memory cap. Full details and complete plots are provided in §6 of [13].
D.3 Summary
[…]
Appendix E Controlled Aliasing: Correctness and Implementation
Preface. Section 5.2 contains a self-contained summary of Controlled Aliasing (including the CNF expansion policy, the $2\times 2$ compensatory merge, and Theorem 6). This appendix provides the full proofs and implementation details.
Navigation. For ease of reading, the lemma names and labels in this appendix match those cited in §5.2: Lemma 8 (state-space reduction), Lemma 6 ($\to10$ shadowing), Lemma 7 (successor preservation), Lemma 9 (expansion accounting), Lemma 10 (policy safety), and the lane/merge lemmas (Lemmas 1113).
To ensure correctness under value aliasing, we introduce a 2-lane memoization scheme keyed by an aliased sum and a 1-bit correction tag. Each reachable aliased sum stores up to two canonical bitmasks, corresponding to whether the substituted element $x_{1}$ is absent or present in the true subset. This enables both decision and construction variants of Subset–Sum to be solved without error, even when synthetic collisions are introduced.
Memoization Structure
Fix an alias pair $(x_{0},x_{1})$ in a half $\ell$ of size $k$ at positions $(i_{0},i_{1})$ with $x_{0}\neq x_{1}$. For a subset $P\subseteq\ell$ with local bitmask $b\in\{0,1\}^{k}$, define the tag
$\chi(P):=\mathbf{1}[\,x_{1}\in P\,]=b[i_{1}],$
and the aliased value
$\tilde{\sigma}(P)\;:=\;\sigma(P)\;-\;\chi(P)\cdot(x_{1}-x_{0}),\qquad\text{so that}\qquad\sigma(P)=\tilde{\sigma}(P)+\chi(P)\cdot(x_{1}-x_{0}).$
We maintain
$\texttt{Memo}:\ \mathbb{Z}\longrightarrow\big{(}\{\bot\}\cup\{0,1\}^{k}\big{)}^{2},\qquad\texttt{Memo}[\tilde{\sigma}][\chi]\in\{\bot\}\cup\{0,1\}^{k},$
storing, for each key $(\tilde{\sigma},\chi)$, the lexicographically minimal witness bitmask (in the local index order of $\ell$).
Canonical Normal Form (CNF).
Define the normalization map $\mathsf{canon}:\{0,1\}^{k}\!\to\!\{0,1\}^{k}$ at the alias indices $(i_{0},i_{1})$ by
$\mathsf{canon}(b)\;=\;\begin{cases}b,&\text{if }b[i_{0}]+b[i_{1}]\in\{0,2\}\ \text{or}\ (b[i_{0}],b[i_{1}])=(1,0),\\ b^{\prime}&\text{if }(b[i_{0}],b[i_{1}])=(0,1),\end{cases}$
where $b^{\prime}$ is $b$ with the pair $(i_{0},i_{1})$ flipped from $(0,1)$ to $(1,0)$. The enumerator enqueues/expands only canonical states $b=\mathsf{canon}(b)$; non-canonical states may still be stored in Memo for correctness but are flagged doNotExtend (non-expandable). This policy is scheduler- and target-independent.
Lemma 6 ($\to10$ shadowing under CNF).
For any bitmask $b$ with $(b[i_{0}],b[i_{1}])=(0,1)$ there exists $b^{\prime}=\mathsf{canon}(b)$ with $(1,0)$ such that $\tilde{\sigma}(b)=\tilde{\sigma}(b^{\prime})$ and $\chi(b)=\chi(b^{\prime})$. Moreover, every aliased extension of $b$ by any $j\notin\{i_{0},i_{1}\}$ is aliased-equal to the corresponding extension of $b^{\prime}$.
Proof.
Normalization replaces $x_{1}$ by $x_{0}$ at $(i_{0},i_{1})$, which preserves $\tilde{\sigma}$ and also $\chi=b[i_{1}]$. Extensions by $j\notin\{i_{0},i_{1}\}$ add the same value to both sides. ∎
Lemma 7 (Successor preservation under CNF).
Let $b$ be any bitmask and $j\notin\{i_{0},i_{1}\}$. Then
$\tilde{\sigma}(b\cup\{j\})=\tilde{\sigma}(\mathsf{canon}(b)\cup\{j\}),\qquad\chi(b\cup\{j\})=\chi(\mathsf{canon}(b)\cup\{j\}).$
Proof.
Immediate from linearity of $\tilde{\sigma}$ and the fact that $\mathsf{canon}$ only possibly flips $(0,1)\!\mapsto\!(1,0)$ at $(i_{0},i_{1})$ with the same aliased contribution. ∎
@@ Appendix E Controlled Aliasing: Correctness and Implementation
Let $S^{\prime}\subseteq\mathbb{Z}_{\geq 0}$ be a set of $k$ non-negative integers, and let $x_{0},x_{1}\in S^{\prime}$ be a designated alias pair with $x_{0}\neq x_{1}$. If $x_{1}$ is treated as $x_{0}$ when present, the number of distinct aliased subset sums is at most $\frac{3}{4}\cdot 2^{k}$.
@@
Fix $P\subseteq S^{\prime}\setminus\{x_{0},x_{1}\}$. The four patterns $P$, $P\cup\{x_{0}\}$, $P\cup\{x_{1}\}$, $P\cup\{x_{0},x_{1}\}$ yield aliased values $\sigma(P)$, $\sigma(P)+x_{0}$, $\sigma(P)+x_{0}$, $\sigma(P)+2x_{0}$, i.e., at most three distinct values. There are $2^{k-2}$ choices of $P$. ∎
Expansion Policy for Worst-Case Bound (count-based).
To tie running time to the number of expanded states (not just distinct aliased values), we restrict which discovered states are enqueued when aliasing is enabled. Let
$c(b):=b[i_{0}]+b[i_{1}]\in\{0,1,2\},\qquad\chi(b):=b[i_{1}]\in\{0,1\}.$
A state with mask $b$ is expandable iff
$\mathrm{Expandable}(b)\;:\iff\;(c(b)=0)\ \lor\ (c(b)=2)\ \lor\ \big{(}c(b)=1\ \wedge\ \chi(b)=0\big{)}.$
Operationally, every discovered state is inserted into its $(\tilde{\sigma},\chi)$ bucket (for correctness and witness recovery), but it is enqueued for expansion iff $\mathrm{Expandable}(b)$ holds. In particular, $(c,\chi)=(1,1)$ ($x_{1}$ only”) states are recorded but marked doNotExtend.
Lemma 9 (Expansion Accounting).
Consider a half $\ell$ of size $k$ with a fixed alias pair $(x_{0},x_{1})$ and the CNF expansion rule. For each base subset $P\subseteq\ell\setminus\{x_{0},x_{1}\}$, the four patterns on $\{x_{0},x_{1}\}$ contribute at most three expanded states. Consequently, the number of expanded states in the half is at most $3\cdot 2^{k-2}=\tfrac{3}{4}\cdot 2^{k}$.
Proof.
Partition subsets by their restriction to $\{i_{0},i_{1}\}$. For each base $P\subseteq\ell\setminus\{x_{0},x_{1}\}$, we expand $00$ and $11$, and among $\{10,01\}$ only $10$. Thus at most three expanded states per base, giving $3\cdot 2^{k-2}$. ∎
Lemma 10 (Policy Safety).
The expansion policy preserves the lane invariant and completeness of the merge: forbidding expansion from $(c,\chi)=(1,1)$ does not prevent the generation of any aliased sum nor the recovery of any witness.
Proof sketch.
For $b$ with $(c,\chi)=(1,1)$, let $b^{\star}$ flip $(i_{0},i_{1})$ to $(1,0)$. For any $J\subseteq\ell\setminus\{x_{0},x_{1}\}$, $\tilde{\sigma}(b\cup J)=\tilde{\sigma}(b^{\star}\cup J)$, so all descendants of $b$ are shadowed by descendants of the expandable $b^{\star}$. $(1,1)$ states are still inserted into the $\chi=1$ bucket, allowing the compensatory merge to return witnesses that include $x_{1}$. ∎
Core Procedures (per half)
Tag and Canonicality.
1:function ChiTag($b$, $i_{1}$)
2:  return $b[i_{1}]$ $\triangleright$ $\chi\in\{0,1\}$
3:function IsCanonical($b$, $i_{0}$, $i_{1}$)
4:  return $b=\mathsf{canon}(b)$
Insert into Memo Table.
1:function Insert($\tilde{\sigma}$, $b$, $i_{1}$)
2:  $\chi\leftarrow\textsc{ChiTag}(b,i_{1})$
@@
4:   Memo[$\tilde{\sigma}$$\chi$] $\leftarrow b$
Expansion Predicate (aliased runs).
1:function Count($b$, $i_{0}$, $i_{1}$)
2:  return $b[i_{0}]+b[i_{1}]$
3:function Expandable($b$, $i_{0}$, $i_{1}$)
4:  $c\leftarrow\textsc{Count}(b,i_{0},i_{1})$, $\chi\leftarrow\textsc{ChiTag}(b,i_{1})$
@@
Hook. When a new state is created under aliasing, set
$\texttt{state.doNotExtend}\;\leftarrow\;\neg\,\textsc{Expandable}(\texttt{state.mask},i_{0},i_{1}).$
Aliasing Maps.
1:function CorrectedSum($\tilde{\sigma}$, $\chi$, $x_{0}$, $x_{1}$)
2:  return $\tilde{\sigma}+\chi\cdot(x_{1}-x_{0})$
3:function AliasedValue($\sigma_{\mathrm{real}}$, $\chi$, $x_{0}$, $x_{1}$)
4:  return $\sigma_{\mathrm{real}}-\chi\cdot(x_{1}-x_{0})$
Interface with CheckAliased.
In the solver (Alg. 6), replace the single-key probe by the compound key $(\tilde{\sigma},\chi)$, computed via AliasedValue and ChiTag. Use IsCanonical/Expandable to gate expansion (CNF). Cross-half tests call the 2-lane CheckAliased routine.
@@
To ensure correctness during merging, the algorithm performs a $2\times 2$ compensatory check across $\chi$-tags of both halves. For each left-side aliased sum, it computes the required right-side partner under both interpretations.
@@
1:function FindSolution($t$, MemoLeft, MemoRight, $x_{0}$, $x_{1}$, $y_{0}$, $y_{1}$)
2:  for all $\tilde{\sigma}_{L}\in\texttt{MemoLeft}$ do
3:   for $\chi_{L}\in\{0,1\}$ do
4:     $b_{L}\leftarrow\texttt{MemoLeft}[\tilde{\sigma}_{L}][\chi_{L}]$
@@
6:     $s_{L}\leftarrow\textsc{CorrectedSum}(\tilde{\sigma}_{L},\chi_{L},x_{0},x_{1})$
@@
8:     for $\chi_{R}\in\{0,1\}$ do
9:      $\tilde{\sigma}_{R}\leftarrow\textsc{AliasedValue}(s_{R},\chi_{R},y_{0},y_{1})$
10:      if $\tilde{\sigma}_{R}\in\texttt{MemoRight}$ then
11:        $b_{R}\leftarrow\texttt{MemoRight}[\tilde{\sigma}_{R}][\chi_{R}]$
@@
13:         return $b_{L}\,\|\,b_{R}$ $\triangleright$ concatenate local masks into the global mask
@@
  • Each reachable aliased sum stores up to two bitmasks (one per $\chi\in\{0,1\}$).
  • All probes are $O(1)$ expected time with hashing; lex-compare occurs only on true collisions.
  • Under the expansion policy, the number of expanded states per half is at most $\tfrac{3}{4}\cdot 2^{k}$ (Lemma 9); this drives the worst-case runtime accounting.
  • Time/space per half: $O(E\cdot n^{2})$ deterministic or $O(E\cdot n)$ randomized, with $E\leq\tfrac{3}{4}\cdot 2^{k}$; storage $O(U^{\prime}\cdot n)$ where $U^{\prime}$ is the aliased-sum count.
Correctness of Controlled Aliasing: Lemma Chain
Setup and notation.
Fix one half $\ell$ (let $k:=|\ell|$) and its alias pair $(x_{0},x_{1})$ at positions $(i_{0},i_{1})$ with $x_{0}\neq x_{1}$. For any subset $P\subseteq\ell$ with local bitmask $b\in\{0,1\}^{k}$, write $\chi(P):=b[i_{1}]\in\{0,1\}$. Let $\sigma(P)$ denote the true sum and $\tilde{\sigma}(P)$ the aliased sum. For a reachable value $\tau$, the table stores a 2-tuple $\mathrm{Memo}[\tau]\in(\{\bot\}\cup\{0,1\}^{k})^{2}$, with the lex-minimal bitmask in bucket $\chi$ if any.
Lemma 11 (Lane Invariant).
For every subset $P\subseteq\ell$, the pair $(\tilde{\sigma}(P),\chi(P))$ identifies a unique bucket in the memoization table. During enumeration, the algorithm maintains for each fixed $(\tau,\chi)$ the lexicographically minimal bitmask among all $P$ with $\tilde{\sigma}(P)=\tau$ and $\chi(P)=\chi$.
Proof.
$\chi$ depends only on the inclusion bit at $i_{1}$, partitioning subsets into two classes. The update rule keeps, for each $(\tilde{\sigma},\chi)$, the lex-minimal witness under the total order $\prec$. ∎
Lemma 12 (Soundness of Compensatory Merge).
If the merge routine returns a witness $W=W_{L}\cup W_{R}$ (with $W_{L}\subseteq\ell_{0}$ and $W_{R}\subseteq\ell_{1}$), then $\sigma(W_{L})+\sigma(W_{R})=t$.
Proof.
The merge checks the four corrected equalities
$\tilde{\sigma}(W_{L})+\tilde{\sigma}(W_{R})+\chi_{R}(y_{1}-y_{0})+\chi_{L}(x_{1}-x_{0})=t$
for all $\chi_{L},\chi_{R}\in\{0,1\}$. As $\sigma(\cdot)=\tilde{\sigma}(\cdot)+\chi(\cdot)\cdot(\cdot)$, any returned witness satisfies $\sigma(W_{L})+\sigma(W_{R})=t$. ∎
Lemma 13 (Completeness of Compensatory Merge).
If there exists a solution $W=W_{L}\cup W_{R}$, then the merge routine will find and return a valid witness.
Proof.
For a valid solution, the two halves appear in buckets $(\tilde{\sigma}(W_{L}),\chi(W_{L}))$ and $(\tilde{\sigma}(W_{R}),\chi(W_{R}))$ by the Lane Invariant. The merge enumerates these tag pairs and restores true sums, so the corresponding check succeeds. ∎
Conclusion.
Lemmas 1113 establish correctness of the 2-lane structure. Together with Lemmas 910, they justify the expansion-based accounting used in Theorem 6. The CNF policy is target-independent and scheduler-agnostic, and, combined with compensatory merging, preserves both decision and construction correctness.
Complexity and Target Independence (Summary)
The Controlled Aliasing transform is purely structural (fixed alias pairs per half), independent of the target $t$ and any randomness. Under CNF, each base pattern contributes to at most three expanded states, yielding the worst-case factor $3/4$ per half and the $O^{*}(2^{\,n/2-\log_{2}(4/3)})$ bound when integrated into the interleaved double–meet-in-the-middle solver.